OpenStudy (anonymous):

find the limit of the function by using direct substitution limits (2e^x cosx)

hartnn (hartnn):

what does x approach ? 0

OpenStudy (anonymous):

pi/2

hartnn (hartnn):

then plug in x=pi/2 in your function \(\large 2e^{\pi/2}\cos (\pi/2)=...?\)

OpenStudy (anonymous):

0?

hartnn (hartnn):

thats correct! :)

OpenStudy (anonymous):

thank you! can you help me on one more?

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

lim x^2+2x/x^4 x approaches to 0

hartnn (hartnn):

is it \(\Large \dfrac{x^2+2x}{x^4}\) ?

OpenStudy (anonymous):

yes, it said to find the limit of the function algebraically

hartnn (hartnn):

have u learnt the concept of left hand and right hand limit ?

OpenStudy (anonymous):

no, not yet

hartnn (hartnn):

that question requires those concepts....

hartnn (hartnn):

u know what x approaches 0, mean ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

what does it mean ?

OpenStudy (anonymous):

isnt it x is approaching to 0 on the right side?

hartnn (hartnn):

for the right side only, we would say \(x \to 0^+\) means x is very near to 0 , but greater than 0 for the left side only, we would say \(x \to 0^-\) means x is very near to 0 , but less than 0 just \(x \to 0\), only means x is very near to 0 see if u get this.

OpenStudy (anonymous):

oh i understand now

OpenStudy (anonymous):

i think i got it, can you check my work? x^2+2x/x^4= x(x+2)/x^4= x+2/x^3= 2/0=does not exist

hartnn (hartnn):

that is when x \(\to 0^+\) we get 2/0= \(+\infty\) but when x \(\to 0^-\) we get -2/0= \(-\infty\) and since these 2 limits are NOT EQUAL, the original limit will NOT EXIST.

OpenStudy (anonymous):

ah i understand, thank you so much!

hartnn (hartnn):

welcome ^_^